OpenStudy (anonymous):

Okay last time I am going to try this problem. Aki's Bicycle Designs has determined that when x hundred bicyc,es are built, the average cost per bicycle is given by C(x)=0.5x^2-0.8x+6.192, where C(x) is in hundreds of dollars. How many bicycles should the shop build to minimize the average cost per bicycle?

OpenStudy (anonymous):

I have worked this problem multiple times and can never get the right answer.

OpenStudy (anonymous):

vertex is \[-\frac{b}{2a}\] with \[a=.5,b=-0.8\]

OpenStudy (anonymous):

so you get \[-\frac{-.8}{2\times .5}=.8\]

OpenStudy (anonymous):

so .8 of 100 is 80. should build 80 bikes.

OpenStudy (anonymous):

thank you